### v3

# Design of Pulse Shapes Based on Sampling with Gaussian Prefilter

Abstract
Two new pulse shapes for communications are presented. The first pulse shape generates a set of pulses without intersymbol interference (ISI) or ISI-free for short. In the neighbourhood of the origin it is similar in shape to the classical cardinal sine function but is of exponential decay at infinity. This pulse shape is identical to the interpolating function of a recent sampling theorem with Gaussian prefilter. The second pulse shape is obtained from the first pulse shape by spectral factorization. Besides being also of exponential decay at infinity, it has a causal appearance since it is of superexponential decay for negative times. It is closely related to the orthonormal generating function considered earlier by Unser in the context of shift-invariant spaces. This pulse shape is not ISI-free but it generates a set of orthonormal pulses. The second pulse shape may also be used to define a receive matched filter so that at the filter output the ISI-free pulses of the first kind are recovered.
I Introduction
Unser [1] extended the standard sampling paradigm to the representation or even approximation of functions by elements of shift-invariant function spaces. These function spaces are defined by a generating function $\varphi$, which has to satisfy certain conditions. By one of them, the partition of unity condition, Gaussian functions are actually precluded from being used as generators. In [2] it was substantiated that Gaussian functions still could be useful generators. Consequently, following [1, Tab. 1], the interpolating generating function and the dual generating function have been computed for a Gaussian generator in [2]. The computation of the corresponding orthonormal generating function, $\varphi_{\mathrm{ortho}}$, is accomplished in Section III of the present paper. Rather than extracting the square root as suggested in [1], our approach is based on spectral factorization and leads, interestingly enough, to expressions in terms of $q$-analogs [3]. As an application, two new pulse shapes are proposed and discussed in Section IV.
The following notations and conventions are adopted: $L^{2}(\mathbb{R})$ is the space of square-integrable functions (or finite-energy signals) $f:\mathbb{R}\rightarrow\mathbb{C}\cup\{\infty\}$ with inner product $\langle f_{1},f_{2}\rangle=\int_{-\infty}^{\infty}f_{1}(x)\overline{f_{2}(x)}\,\mathrm{d}x.$ For the Fourier transform we use the definition $\hat{f}(\omega)=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}f(x)\,\mathrm{d}x$, where $x$ denotes time and $\omega$ angular frequency. $\ell^{2}(\mathbb{Z})$ is the space of square-summable, complex-valued sequences indexed by the integers. Finally, $\delta_{n}=0,\,n\in\mathbb{Z}\setminus\{0\}$, and $\delta_{0}=1$.
II Sampling in Shift-Invariant Spaces and Localization Spaces Revisited
The purpose of the present section is to motivate the pulse shapes presented in Section IV. To this end, we give a brief overview of sampling in shift-invariant spaces and localization spaces with an emphasis on those spaces defined by a Gaussian generator or a Gaussian prefilter respectively.
Suppose that $\varphi\in L^{2}(\mathbb{R})$ is a continuous function satisfying $\varphi(x)=O(|x|^{-1-\epsilon})$ as $\,x\rightarrow\pm\infty$ for some $\epsilon>0$ and for any $\lambda>0$ the system of functions $\{\varphi(\cdot-n\lambda);n\in\mathbb{Z}\}$ forms a Riesz basis in $L^{2}(\mathbb{R})$. Furthermore, suppose that for any $\lambda>0$
$\sum_{n=-\infty}^{\infty}\varphi(n\lambda)\mathrm{e}^{-\mathrm{i}n\lambda\omega}\not=0,\,\omega\in\mathbb{R}.$ (1)
The shift-invariant space $V_{\lambda}(\varphi)$ is the subspace of $L^{2}(\mathbb{R})$ defined as
$V_{\lambda}(\varphi)=\left\{f;f(x)=\sum_{n\in\mathbb{Z}}c_{n}\varphi(x-n\lambda),\,c\in\ell^{2}(\mathbb{Z})\right\}.$ (2)
Then, the following sampling theorem applies.
Theorem 1
For any $\lambda>0$ and $f\in V_{\lambda}(\varphi)$ it holds that
$f(x)=\sum_{n\in\mathbb{Z}}f(n\lambda)\varphi_{\mathrm{int}}(x-n\lambda),\,x\in\mathbb{R},$ (3)
where the interpolating function $\varphi_{\mathrm{int}}\in V_{\lambda}(\varphi)$ is given by
$\hat{\varphi}_{\mathrm{int}}(\omega)$ $=\frac{\hat{\varphi}(\omega)}{\sum_{n\in\mathbb{Z}}\varphi(n\lambda)\mathrm{e}^{-\mathrm{i}n\lambda\omega}}$
$=\frac{\hat{\varphi}(\omega)}{\frac{\Lambda}{\sqrt{2\pi}}\sum_{k\in{\mathbb{Z}}}\hat{\varphi}(\omega+k\Lambda)},\quad\Lambda=\frac{2\pi}{\lambda}.$ (4)
This theorem is due to Walter [4], who originally stated and proved it for orthonormal systems $\{\varphi(\cdot-n);n\in\mathbb{Z}\}$. The theorem was later generalized to Riesz bases by Unser [1], who also considered alternative generating functions for $V_{\lambda}(\varphi)$ like $\varphi_{\mathrm{int}}$ as above and $\varphi_{\mathrm{ortho}}$ (see below). In an attempt to retain the flavor of the original Whittaker–Kotelnikov–Shannon (WKS) sampling theorem [5], in [6] prior to sampling a prefilter
$(\boldsymbol{P}_{\varphi}f)(x)=\int_{-\infty}^{\infty}f(y)\overline{\varphi(y-x)}\,\mathrm{d}y$ (5)
with prefilter function $\varphi\in L^{2}(\mathbb{R})$ has been applied to an arbitrary finite-energy signal $f\in L^{2}(\mathbb{R})$. The so-called localization space
$\mathcal{P}_{\varphi}=\{g=\boldsymbol{P}_{\varphi}f;f\in L^{2}({\mathbb{R}})\}$
then corresponds to the space of bandlimited, finite-energy signals in the classical WKS sampling theorem. The goal is to recover the filter output signal $g=\boldsymbol{P}_{\varphi}f$ from sample values $g(n\lambda),\,n\in\mathbb{Z},$ either perfectly or at least with an acceptable error. To this end, the autocorrelation function $\Phi$ of $\varphi$,
$\Phi=\boldsymbol{P}_{\varphi}\varphi\in\mathcal{P}_{\varphi}\stackrel{{\scriptstyle\mathrm{Fourier}}}{{\longleftrightarrow}}\hat{\Phi}(\omega)=\sqrt{2\pi}|\hat{\varphi}(\omega)|^{2},$ (6)
is needed. The (second) interpolating function $\Phi_{\mathrm{int}}\in\mathcal{P}_{\varphi}$ is defined by its Fourier transform
$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\hat{\Phi}(\omega)}{\frac{\Lambda}{\sqrt{2\pi}}\sum_{k\in{\mathbb{Z}}}\hat{\Phi}(\omega+k\Lambda)},$ (7)
where $\Lambda$ is as in (4). Note that because of the Riesz basis condition still imposed on $\varphi$ (see [6] for the full set of assumptions), which is equivalent to the existence of positive constants $A$ and $B$ (possibly depending on $\lambda$) so that
$0<A\leq\Lambda\!\sum_{k=-\infty}^{\infty}|\hat{\varphi}(\omega+k\Lambda)|^{2}\leq B<\infty,\;\;\omega\in\mathbb{R},$ (8)
(see [7] for a proof in the case of $\lambda=1$), the denominator in (7) never will vanish. We remark that in general $\Phi_{\mathrm{int}}\neq\boldsymbol{P}_{\varphi}\varphi_{\mathrm{int}}$. In the special case of a Gaussian prefilter function,
$\varphi(x)\!=\!\frac{1}{\sqrt{2\pi}(1/\beta)}\,\mathrm{e}^{-\frac{x^{2}}{2(1/\beta)^{2}}}\stackrel{{\scriptstyle\mathrm{Fourier}}}{{\longleftrightarrow}}\hat{\varphi}(\omega)\!=\!\frac{1}{\sqrt{2\pi}}\,\mathrm{e}^{-\frac{\omega^{2}}{2\beta^{2}}},$ (9)
where the parameter $\beta>0$ controls effective bandwidth, the following sampling theorem has been obtained [2], [6].
Theorem 2
For any $\lambda>0$ let the interpolating function $\Phi_{\mathrm{int}}$ be defined by (7). Then for any function $g\in\mathcal{P}_{\varphi}$ the function $\tilde{g}$ given by
$\tilde{g}(x)=\sum_{n\in\mathbb{Z}}g(n\lambda)\Phi_{\mathrm{int}}(x-n\lambda),\;x\in\mathbb{R},$ (10)
is again in $\mathcal{P}_{\varphi}$, it perfectly reconstructs $g$ at the sampling instants $x_{n}=n\lambda,\,n\in\mathbb{Z}$, and for all other $x\in\mathbb{R}$ the absolute approximation error $|g(x)-\tilde{g}(x)|$ becomes small as soon as $\lambda\leq 1/\beta$, then decaying superexponentially to zero as $\lambda\rightarrow 0$.
For the Gaussian prefilter function $\varphi$ as given in (9) (which will be assumed for the rest of the paper) one has
$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\lambda}{\sqrt{2\pi}}\frac{\mathrm{e}^{-\frac{i\pi}{\tau}(\omega/\Lambda)^{2}}}{\sqrt{-i\tau}\vartheta_{3}(\omega/\Lambda,\tau)},$ (11)
where
$\vartheta_{3}(z,\tau)=\frac{1}{\sqrt{-\mathrm{i}\tau}}\sum_{n=-\infty}^{\infty}\mathrm{e}^{-\frac{\mathrm{i}\pi}{\tau}(z+n)^{2}}$ (12)
is a Jacobi theta function [8] with parameter $\tau$ given by
$\tau=\mathrm{i}(\lambda\beta)^{2}\!/(4\pi).$ (13)
By inversion of the Fourier transform we obtain after use of Jacobi’s $\tau\rightarrow-1/\tau$ transformation that [2], [9]
$\Phi_{\mathrm{int}}(x)=\frac{\mathrm{i}\pi\tau}{\vartheta_{1}^{\prime}\left(0,-1/\tau\right)}\frac{\vartheta_{1}\left(x/\lambda,-1/\tau\right)}{\sinh(\mathrm{i}\pi\tau x/\lambda)},$ (14)
where $\vartheta_{1}(z,\tau)=2\sum_{n=0}^{\infty}q^{\left(n+\frac{1}{2}\right)^{2}}(-1)^{n}\sin[(2n+1)\pi z],\,q=\mathrm{e}^{\mathrm{i}\pi\tau},\,\Im(\tau)>0,$ is another theta function [8]. When $\lambda\leq 1/\beta$, it holds with high accuracy that $\Phi_{\mathrm{int}}(x)\approx S_{0}(x)$, where [9]
$S_{0}(x)\triangleq\mathrm{i}\tau\frac{\sin(\pi x/\lambda)}{\sinh(\mathrm{i}\pi\tau x/\lambda)},\;x\in\mathbb{R}.$ (15)
Note that in our context $\mathrm{i}\tau$ is always a negative real number and that $S_{0}(x)$ decays exponentially to zero as $x\rightarrow\pm\infty$.
Fig. 1 shows the interpolating function $\Phi_{\mathrm{int}}$ for $\beta=100$ and $\lambda=1/\beta$; actually, the approximation (15) for $\Phi_{\mathrm{int}}(x)$ has been used.
III Spectral Factorization of the Interpolating Function
We start by compiling a few prerequisites [3].
Definition 1
For any $a\in{\mathbb{R}}$ and $q\in{\mathbb{C}}$ with $|q|<1$ the $q$-Pochhammer symbol $(a;q)_{n}$ is defined by
$(a;q)_{n}$ $=(1-a)(1-aq)\cdots(1-aq^{n-1}),\,n=1,2,\ldots,$
$(a;q)_{\infty}$ $=\lim_{n\rightarrow\infty}(a;q)_{n}\,,$
$(a;q)_{0}$ $=1.$
The following identity of Euler holds true for any $q\in{\mathbb{C}},|q|<1,$ and $z\in\mathbb{C}$:
$1+\sum_{n=1}^{\infty}\frac{q^{\frac{1}{2}n(n-1)}}{(q;q)_{n}}z^{n}=\prod_{n=0}^{\infty}(1+zq^{n}).$ (16)
Jacobi’s triple product identity
$\sum_{n=-\infty}^{\infty}q^{n^{2}}z^{n}=\prod_{n=1}^{\infty}(1-q^{2n})(1+q^{2n-1}z^{-1})(1+q^{2n-1}z)$ (17)
will also be needed. In our case always
$q=\mathrm{e}^{\mathrm{i}\pi\tau},$ (18)
where the parameter $\tau$ is as in (13). The special $q$-Pochhammer symbols
$(q^{2};q^{2})_{n}=\prod_{k=1}^{n}(1-q^{2k}),\quad Q_{0}\triangleq(q^{2};q^{2})_{\infty}$
will occur frequently.
Theorem 3
For the function
$\varphi_{\mathrm{ortho}}(x)=\frac{1}{\sqrt{(q^{2};q^{2})_{\infty}}}\sum_{n=0}^{\infty}\frac{(-1)^{n}q^{n}}{(q^{2};q^{2})_{n}}\,\phi(x-n\lambda),$ (19)
where
$\phi(x)=\frac{\beta^{1/2}}{\pi^{1/4}}\,\mathrm{e}^{-\frac{x^{2}}{2(1/\beta)^{2}}}$ (20)
is the Gaussian function (9) normalized to unit energy, i.e. $\int_{\mathbb{R}}|\phi(x)|^{2}\,\mathrm{d}x=1$, it holds that
$\hat{\Phi}_{\mathrm{int}}(\omega)=\sqrt{2\pi}|\hat{\varphi}_{\mathrm{ortho}}(\omega)|^{2},\;\omega\in\mathbb{R}.$ (21)
Proof:
The definition (11) of the interpolating function $\Phi_{\mathrm{int}}$ may be written in the Fourier domain as
$\hat{\Phi}_{\mathrm{int}}(\omega)=\sqrt{2\pi}\lambda\frac{|\hat{\varphi}(\omega)|^{2}}{\sqrt{-\mathrm{i}\tau}\vartheta_{3}(\omega/\Lambda,\tau)}.$
Then, Eq. (21) becomes
$|\hat{\varphi}_{\mathrm{ortho}}(\omega)|^{2}=\lambda\frac{|\hat{\varphi}(\omega)|^{2}}{\sqrt{-\mathrm{i}\tau}\vartheta_{3}(\omega/\Lambda,\tau)}.$
Since the theta function (12) has the second representation [8]
$\vartheta_{3}(x,\tau)=1+2\sum_{n=0}^{\infty}q^{n^{2}}\cos(2\pi nx)=\sum_{n=-\infty}^{\infty}q^{n^{2}}\mathrm{e}^{2\pi\mathrm{i}nx},$
we obtain by means of Eq. (17), putting
$P(z)=\prod_{n=1}^{\infty}(1+q^{2n-1}z^{-1})$
and subsequently $z=\mathrm{e}^{2\pi\mathrm{i}x}$, for $\vartheta_{3}(x,\tau)$ the factorization
$\vartheta_{3}(x,\tau)=Q_{0}^{1/2}P(\mathrm{e}^{2\pi\mathrm{i}x})\cdot\overline{Q_{0}^{1/2}P(\mathrm{e}^{2\pi\mathrm{i}x})},\;x\in\mathbb{R}.$
Since the function $x\mapsto\vartheta_{3}(x,\tau)$ is real-valued and positively lower bounded on $\mathbb{R}$ so is the function $x\mapsto|P(\mathrm{e}^{2\pi\mathrm{i}x})|$. As a consequence, the definition of $\varphi_{\mathrm{ortho}}$ in the Fourier domain by
$\hat{\varphi}_{\mathrm{ortho}}(\omega)=\lambda^{\frac{1}{2}}\frac{\hat{\varphi}(\omega)}{(-\mathrm{i}\tau)^{1/4}Q_{0}^{1/2}P(\mathrm{e}^{2\pi i\omega/\Lambda})}$
will result in a function $\varphi_{\mathrm{ortho}}\in L^{2}(\mathbb{R})$ satisfying Eq. (21).
Now, we need to invert the Fourier transform. Since $x\mapsto 1/P(\mathrm{e}^{2\pi\mathrm{i}x})$ is a bounded 1-periodic function, it is in $L^{2}([0,1))$ and thus has a Fourier series expansion $1/P(\mathrm{e}^{2\pi\mathrm{i}x})=\sum_{n=-\infty}^{\infty}a_{n}\mathrm{e}^{-2\pi\mathrm{i}nx},\,a\in\ell^{2}(\mathbb{Z}),$ with the coefficients
$a_{n}=\int_{0}^{1}\mathrm{e}^{2\pi\mathrm{i}nx}\frac{1}{P(\mathrm{e}^{2\pi\mathrm{i}x})}\,\mathrm{d}x,\,n\in\mathbb{Z}.$ (22)
By inverse Fourier transform we now obtain (putting $c=\lambda^{1/2}(-\mathrm{i}\tau)^{-1/4}Q_{0}^{-1/2}$) that
$\varphi_{\mathrm{ortho}}(x)$ $=$ $\frac{c}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}x\omega}\frac{\hat{\varphi}(\omega)}{P(\mathrm{e}^{2\pi\mathrm{i}\omega/\Lambda})}\,\mathrm{d}\omega$
$=$ $c\sum_{n=-\infty}^{\infty}\frac{a_{n}}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}(x-n\lambda)\omega}\hat{\varphi}(\omega)\,\mathrm{d}\omega$
$=$ $c\sum_{n=-\infty}^{\infty}a_{n}\varphi(x-n\lambda)$
$=$ $\frac{1}{Q_{0}^{1/2}}\sum_{n=-\infty}^{\infty}a_{n}\phi(x-n\lambda).$
The computation of the coefficients (22) is carried out in the complex domain.
Case $n=-1,-2,\ldots:$ After substitution in (22) of $x$ by $1-x$ we obtain
$a_{n}$ $=$ $\int_{0}^{1}\mathrm{e}^{-2\pi\mathrm{i}nx}\frac{1}{P(\mathrm{e}^{-2\pi\mathrm{i}x})}\,\mathrm{d}x$
$=$ $\frac{1}{2\pi\mathrm{i}}\oint_{|z|=1}\frac{z^{-n-1}}{P(z^{-1})}\,\mathrm{d}z,$
where integration in the contour integral is performed counterclockwise around the unit circle. Since the integrand function is analytic within a neighbourhood of the closed unit disc, we obtain by means of Cauchy’s integral theorem that $a_{n}=0,\,n=-1,-2,\ldots$
Case $n=0,1,\ldots:$ Eq. (22) now directly yields by transition to a contour integral that
$a_{n}$ $=$ $\frac{1}{2\pi\mathrm{i}}\oint_{|z|=1}\frac{z^{n-1}}{P(z)}\,\mathrm{d}z$
$=$ $\lim_{M\rightarrow\infty}\underbrace{\frac{1}{2\pi\mathrm{i}}\oint_{|z|=1}\frac{z^{n-1}}{P_{M}(z)}\,\mathrm{d}z}_{I_{M}(n)},$
where
$P_{M}(z)=\prod_{k=0}^{M-1}(1+q^{2k+1}z^{-1}),\,M=1,2,\ldots,$
and the path of integration is same as before. The integrand function in the integral defining $I_{M}(n)$ has simple poles at
$z_{m}=-q^{2m+1},\,m=0,1,\ldots,M-1,$
lying inside of the unit circle. By means of the theorem of residues we obtain
$I_{M}(n)=\sum_{m=0}^{M-1}\mathrm{Res}_{z_{m}}\frac{z^{n-1}}{P_{M}(z)},$
whence
$a_{n}=\lim_{M\rightarrow\infty}I_{M}(n)=\sum_{m=0}^{\infty}\mathrm{Res}_{z_{m}}\frac{z^{n-1}}{P(z)}.$
We compute that
$\mathrm{Res}_{z_{m}}\frac{z^{n-1}}{P(z)}$ $=$ $\mathrm{Res}_{z_{m}}\frac{z^{n}}{zP(z)}$
$=$ $\mathrm{Res}_{z_{m}}\frac{z^{n}}{(z-z_{m})\prod_{k=0,k\neq m}^{\infty}(1+q^{2k+1}z^{-1})}$
$=$ $\frac{z_{m}^{n}}{\prod_{k=0,k\neq m}^{\infty}(1+q^{2k+1}z_{m}^{-1})}$
$=$ $\frac{(-q)^{n}q^{2mn}}{(q^{2};q^{2})_{\infty}\prod_{k=1}^{m}(1-q^{-2k})}$
$=$ $\frac{(-q)^{n}}{(q^{2};q^{2})_{\infty}}\,\frac{(-1)^{m}q^{m(m+1)}(q^{2n})^{m}}{\prod_{k=1}^{m}(1-q^{2k})},$
treating in the case of $m=0$ the empty product as one. After replacement of $q$ in Eq. (16) with $q^{2}$ we get
$1+\sum_{m=1}^{\infty}\frac{q^{m(m-1)}}{(q^{2};q^{2})_{m}}z^{m}=\prod_{m=0}^{\infty}(1+zq^{2m}).$
Hence
$a_{n}$ $=$ $\frac{(-q)^{n}}{(q^{2};q^{2})_{\infty}}\sum_{m=0}^{\infty}\frac{q^{m(m-1)}}{(q^{2};q^{2})_{m}}(-q^{2(n+1)})^{m}$
$=$ $\frac{(-q)^{n}}{(q^{2};q^{2})_{\infty}}\prod_{m=0}^{\infty}(1-q^{2(n+1)}q^{2m})$
$=$ $\frac{(-q)^{n}}{(q^{2};q^{2})_{n}},$
which concludes the proof of Theorem 3. ∎
It can be shown that $V_{\lambda}(\varphi_{\mathrm{ortho}})=V_{\lambda}(\varphi)$, where $V_{\lambda}(\varphi)$ is again the shift-invariant space (2) with Gaussian generating function $\varphi$ as given in (9). In Fig. 2, the function $\varphi_{\mathrm{ortho}}$ together with some of its translates is depicted for bandwidth parameter $\beta=100$ and $\lambda=3/\beta$.
IV Applications
Figure 1: $\Phi_{\mathrm{int}}$ for $\beta=100$ and $\lambda=1/\beta$; sinc is the cardinal sine function defined by $\mathrm{sinc}(x)=\sin(\pi x)/(\pi x)$.
Figure 2: $\varphi_{\mathrm{ortho}}$ together with some of its orthogonal translates for $\beta=100$ and $\lambda=3/\beta$.
IV-A ISI-Free Pulses
The function $\Phi_{\mathrm{int}}$ in (14) may be used as a pulse shape to generate ISI-free pulses. Indeed, from representation (7) it is readily seen that for all $\omega\in\mathbb{R}$ it holds
$\sum_{n\in\mathbb{Z}}\Phi_{\mathrm{int}}(n\lambda)\mathrm{e}^{-\mathrm{i}n\lambda\omega}=\frac{\Lambda}{\sqrt{2\pi}}\sum_{k\in\mathbb{Z}}\hat{\Phi}_{\mathrm{int}}(\omega+k\Lambda)=1$ (23)
(the left-hand equation being Poisson’s summation formula; e.g., [1]), which is equivalent to $\Phi_{\mathrm{int}}(n\lambda)=\delta_{n},\,n\in\mathbb{Z}$. Therefore, the set of shifted pulses $\{\Phi_{\mathrm{int}}(x-n\lambda);n\in\mathbb{Z}\}$ is ISI-free at points in time $x_{n}=n\lambda$. The right-hand equation in (23) is, of course, the well-known Nyquist criterion for ISI-free pulses [10]. Concerning the use of the also ISI-free pulses generated by (15) (for arbitrary $\beta>0,\,\lambda>0$) refer to [11].
IV-B Orthonormal Pulses with ISI-Free Matched Filter Output
IV-B1 Orthonormal Pulses
It is well-known [1] that for an arbitrary function $\varphi\in L^{2}(\mathbb{R})$ the system of functions $\{\varphi(x-n\lambda);n\in\mathbb{Z}\}$ forms an orthonomal system in $L^{2}(\mathbb{R})$ if and only if in (8) $A=B=1$ may be chosen. For the function $\varphi_{\mathrm{ortho}}$ of Theorem 3 this is true due to Eq. (21) and the right-hand equation in (23). Therefore, $\varphi_{\mathrm{ortho}}$ may be used as a pulse shape to generate a set $\{\varphi_{\mathrm{ortho}}(x-n\lambda);\,n\in\mathbb{Z}\}$ of orthonormal pulses.
IV-B2 Matched Filter
At the receiver, the filter $\boldsymbol{P}_{\varphi_{\mathrm{ortho}}}$ obtained by replacing the prefilter function $\varphi$ in (5) with $\varphi_{\mathrm{ortho}}$ forms a matched filter allowing optimal detection of the pulses $\varphi_{\mathrm{ortho}}(x-n\lambda)$ at points in time $x_{n}=n\lambda,\;n\in\mathbb{Z},$ in the presence of noise. Thanks to orthogonality, overlap of consecutive pulses is irrelevant. Moreover, since [cf. (6) and Eq. (21)]
$\boldsymbol{P}_{\varphi_{\mathrm{ortho}}}\varphi_{\mathrm{ortho}}=\Phi_{\mathrm{int}},$
the ISI-free pulses $\Phi_{\mathrm{int}}(x-n\lambda),\,n\in\mathbb{Z},$ of Section IV-A are recovered at the matched filter output.
In the light of the last application, the two proposed pulse shapes $\Phi_{\mathrm{int}}$ and $\varphi_{\mathrm{ortho}}$ are seen to correspond to the classical pulse shapes produced by a raised-cosine filter or a root raised-cosine filter respectively [10]. Recall that the proposed pulse shapes decay (super-)exponentially as $x\rightarrow\pm\infty$ whereas the classical pulse shapes merely decay as $x^{-3}$.
acknowledgement
The author wishes to thank Ayush Bhandari for interesting discussions and helpful suggestions concerning some topics of this paper.

### v4

@@
 The first pulse shape generates a set of pulses without intersymbol interference (ISI) or ISI-free for short.
-In the neighbourhood of the origin it is similar in shape to the classical cardinal sine function but is of exponential decay at infinity.
-This pulse shape is identical to the interpolating function of a recent sampling theorem with Gaussian prefilter.
+In the neighborhood of the origin it is similar in shape to the classical cardinal sine function but is of exponential decay at infinity.
+This pulse shape is identical to the interpolating function of a generalized sampling theorem with Gaussian prefilter.
 The second pulse shape is obtained from the first pulse shape by spectral factorization.
@@
 As an application, two new pulse shapes are proposed and discussed in Section IV.
-The following notations and conventions are adopted: $L^{2}(\mathbb{R})$ is the space of square-integrable functions (or finite-energy signals) $f:\mathbb{R}\rightarrow\mathbb{C}\cup\{\infty\}$ with inner product $\langle f_{1},f_{2}\rangle=\int_{-\infty}^{\infty}f_{1}(x)\overline{f_{2}(x)}\,\mathrm{d}x.$ For the Fourier transform we use the definition $\hat{f}(\omega)=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}f(x)\,\mathrm{d}x$, where $x$ denotes time and $\omega$ angular frequency.
+The following notations and conventions are adopted: $L^{2}(\mathbb{R})$ is the space of square-integrable functions (or finite-energy signals) $f:\mathbb{R}\rightarrow\mathbb{C}\cup\{\infty\}$ with inner product $\langle f_{1},f_{2}\rangle=\int_{-\infty}^{\infty}f_{1}(x)\overline{f_{2}(x)}\,\mathrm{d}x$ and norm $\|f\|=\langle f,f\rangle^{1/2}$.
+For the Fourier transform we use the definition $\hat{f}(\omega)=(2\pi)^{-1/2}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}f(x)\,\mathrm{d}x$, where $x$ denotes time and $\omega$ angular frequency.
 $\ell^{2}(\mathbb{Z})$ is the space of square-summable, complex-valued sequences indexed by the integers.
@@
 $=\frac{\hat{\varphi}(\omega)}{\frac{\Lambda}{\sqrt{2\pi}}\sum_{k\in\mathbb{Z}}\hat{\varphi}(\omega+k\Lambda)},\quad\Lambda=\frac{2\pi}{\lambda}.$ (4)
-This theorem is due to Walter [4], who originally stated and proved it for orthonormal systems $\{\varphi(\cdot-n);n\in\mathbb{Z}\}$.
-The theorem was later generalized to Riesz bases by Unser [1], who also considered alternative generating functions for $V_{\lambda}(\varphi)$ like $\varphi_{\mathrm{int}}$ as above and $\varphi_{\mathrm{ortho}}$ (see below).
+This theorem is due to Walter [4], who originally proved it for orthonormal bases $\{\varphi(\cdot-n);n\in\mathbb{Z}\}$.
+The theorem was later extended to Riesz bases by Unser [1], who also considered alternative generating functions for $V_{\lambda}(\varphi)$ like $\varphi_{\mathrm{int}}$ as above and $\varphi_{\mathrm{ortho}}$ (see below).
 In an attempt to retain the flavor of the original Whittaker–Kotelnikov–Shannon (WKS) sampling theorem [5], in [6] prior to sampling a prefilter
@@
 where $\Lambda$ is as in (4).
-Note that because of the Riesz basis condition still imposed on $\varphi$ (see [6] for the full set of assumptions), which is equivalent to the existence of positive constants $A$ and $B$ (possibly depending on $\lambda$) so that
+Note that because of the Riesz basis condition still imposed on $\varphi$ (see [6] for the full set of assumptions), which is equivalent to the existence of positive constants $A$ and $B$ (possibly depending on $\lambda$) so that [7]
 $0<A\leq\Lambda\!\sum_{k=-\infty}^{\infty}|\hat{\varphi}(\omega+k\Lambda)|^{2}\leq B<\infty,\;\omega\in\mathbb{R},$ (8)
-(see [7] for a proof in the case of $\lambda=1$), the denominator in (7) never will vanish.
-We remark that in general $\Phi_{\mathrm{int}}\neq\boldsymbol{P}_{\varphi}\varphi_{\mathrm{int}}$.
-In the special case of a Gaussian prefilter function,
+the denominator in (7) never will vanish. We remark that in general $\Phi_{\mathrm{int}}\neq\boldsymbol{P}_{\varphi}\varphi_{\mathrm{int}}$. In the special case of a Gaussian prefilter function,
 $\varphi(x)\!=\!\frac{1}{\sqrt{2\pi}(1/\beta)}\,\mathrm{e}^{-\frac{x^{2}}{2(1/\beta)^{2}}}\stackrel{{\scriptstyle\mathrm{Fourier}}}{{\longleftrightarrow}}\hat{\varphi}(\omega)\!=\!\frac{1}{\sqrt{2\pi}}\,\mathrm{e}^{-\frac{\omega^{2}}{2\beta^{2}}},$ (9)
-where the parameter $\beta>0$ controls effective bandwidth, the following sampling theorem has been obtained [2], [6].
+where the parameter $\beta>0$ controls effective bandwidth, the following generalized sampling theorem has been obtained [2], [6].
 Theorem 2
-For any $\lambda>0$ let the interpolating function $\Phi_{\mathrm{int}}$ be defined by (7). Then for any function $g\in\mathcal{P}_{\varphi}$ the function $\tilde{g}$ given by
+For any $\lambda>0$ let the interpolating function $\Phi_{\mathrm{int}}$ be defined by (7).
+Then for any function $g\in\mathcal{P}_{\varphi}$, where $g=\boldsymbol{P}_{\varphi}f,\,f\in L^{2}(\mathbb{R})$, the function $\tilde{g}$ given by
 $\tilde{g}(x)=\sum_{n\in\mathbb{Z}}g(n\lambda)\Phi_{\mathrm{int}}(x-n\lambda),\;x\in\mathbb{R},$ (10)
-is again in $\mathcal{P}_{\varphi}$, it perfectly reconstructs $g$ at the sampling instants $x_{n}=n\lambda,\,n\in\mathbb{Z}$, and for all other $x\in\mathbb{R}$ the absolute approximation error $|g(x)-\tilde{g}(x)|$ becomes small as soon as $\lambda\leq 1/\beta$, then decaying superexponentially to zero as $\lambda\rightarrow 0$.
+is again in $\mathcal{P}_{\varphi}$, it perfectly reconstructs $g$ at the sampling instants $x_{n}=n\lambda,\,n\in\mathbb{Z}$, and for all other $x\in\mathbb{R}$ the squared relative error $\left(|g(x)-\tilde{g}(x)|/\|f\|\right)^{2}$ becomes small as soon as $\lambda\leq 1/\beta$, then decaying superexponentially to zero as $\lambda\rightarrow 0$.
 For the Gaussian prefilter function $\varphi$ as given in (9) (which will be assumed for the rest of the paper) one has
-$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\lambda}{\sqrt{2\pi}}\frac{\mathrm{e}^{-\frac{i\pi}{\tau}(\omega/\Lambda)^{2}}}{\sqrt{-i\tau}\vartheta_{3}(\omega/\Lambda,\tau)},$ (11)
+$\hat{\Phi}_{\mathrm{int}}(\omega)=\frac{\lambda}{\sqrt{2\pi}}\frac{\mathrm{e}^{-\frac{\mathrm{i}\pi}{\tau}(\omega/\Lambda)^{2}}}{\sqrt{-\mathrm{i}\tau}\vartheta_{3}(\omega/\Lambda,\tau)},$ (11)
 where
@@
 $S_{0}(x)\triangleq\mathrm{i}\tau\frac{\sin(\pi x/\lambda)}{\sinh(\mathrm{i}\pi\tau x/\lambda)},\;x\in\mathbb{R}.$ (15)
-Note that in our context $\mathrm{i}\tau$ is always a negative real number and that $S_{0}(x)$ decays exponentially to zero as $x\rightarrow\pm\infty$.
+Note that in our context $\mathrm{i}\tau$ is always a negative real number and that $\Phi_{\mathrm{int}}(x)$ decays exponentially to zero as $x\rightarrow\pm\infty$.
 Fig. 1 shows the interpolating function $\Phi_{\mathrm{int}}$ for $\beta=100$ and $\lambda=1/\beta$; actually, the approximation (15) for $\Phi_{\mathrm{int}}(x)$ has been used.
@@
 $(a;q)_{0}$ $=1.$
-The following identity of Euler holds true for any $q\in{\mathbb{C}},|q|<1,$ and $z\in\mathbb{C}$:
-$1+\sum_{n=1}^{\infty}\frac{q^{\frac{1}{2}n(n-1)}}{(q;q)_{n}}z^{n}=\prod_{n=0}^{\infty}(1+zq^{n}).$ (16)
-Jacobi’s triple product identity
+The following identities of Euler hold true for $q\in\mathbb{C},|q|<1$:
+$1+\sum_{n=1}^{\infty}\frac{q^{\frac{1}{2}n(n-1)}}{(q;q)_{n}}z^{n}=\prod_{n=0}^{\infty}(1+zq^{n}),\,z\in\mathbb{C},$ (16)
+$1+\sum_{n=1}^{\infty}\frac{z^{n}}{(q;q)_{n}}=\prod_{n=0}^{\infty}(1-zq^{n})^{-1},\,z\in\mathbb{C},|z|<1.$ (17)
+Only Eq. (16) will be needed in the proof of the next theorem, where Jacobi’s triple product identity
 $\sum_{n=-\infty}^{\infty}q^{n^{2}}z^{n}=\prod_{n=1}^{\infty}(1-q^{2n})(1+q^{2n-1}z^{-1})(1+q^{2n-1}z)$ (18)
-will also be needed. In our case always
+will play a central role. In our case always
 $q=\mathrm{e}^{\mathrm{i}\pi\tau},$ (19)
-where the parameter $\tau$ is as in (13). The special $q$-Pochhammer symbols
+where parameter $\tau$ is as in (13). The special $q$-Pochhammer symbols
 $(q^{2};q^{2})_{n}=\prod_{k=1}^{n}(1-q^{2k}),\quad Q_{0}\triangleq(q^{2};q^{2})_{\infty}$
@@
 $\varphi_{\mathrm{ortho}}(x)=\frac{1}{\sqrt{(q^{2};q^{2})_{\infty}}}\sum_{n=0}^{\infty}\frac{(-1)^{n}q^{n}}{(q^{2};q^{2})_{n}}\,\phi(x-n\lambda),$ (20)
-where
+where $x\in\mathbb{R}$ and
 $\phi(x)=\frac{\beta^{1/2}}{\pi^{1/4}}\,\mathrm{e}^{-\frac{x^{2}}{2(1/\beta)^{2}}}$ (21)
-is the Gaussian function (9) normalized to unit energy, i.e. $\int_{\mathbb{R}}|\phi(x)|^{2}\,\mathrm{d}x=1$, it holds that
+is the Gaussian function $\varphi$ in (9) normalized to unit energy, i.e., $\int_{\mathbb{R}}|\phi(x)|^{2}\,\mathrm{d}x=1$, it holds that
 $\hat{\Phi}_{\mathrm{int}}(\omega)=\sqrt{2\pi}|\hat{\varphi}_{\mathrm{ortho}}(\omega)|^{2},\;\omega\in\mathbb{R}.$ (22)
@@
 The computation of the coefficients (23) is carried out in the complex domain.
-Case $n=-1,-2,\ldots:$ After substitution in (22) of $x$ by $1-x$ we obtain
+Case $n=-1,-2,\ldots:$ After substitution in (23) of $x$ by 1x1-x we obtain
 $a_{n}$ $=$ $\int_{0}^{1}\mathrm{e}^{-2\pi\mathrm{i}nx}\frac{1}{P(\mathrm{e}^{-2\pi\mathrm{i}x})}\,\mathrm{d}x$
@@
 where integration in the contour integral is performed counterclockwise around the unit circle.
-Since the integrand function is analytic within a neighbourhood of the closed unit disc, we obtain by means of Cauchy’s integral theorem that $a_{n}=0,\,n=-1,-2,\ldots$
+Since the integrand function is analytic within a neighborhood of the closed unit disc, we obtain by means of Cauchy’s integral theorem that $a_{n}=0,\,n=-1,-2,\ldots$
 Case $n=0,1,\ldots:$ Eq. (23) now directly yields by transition to a contour integral that
@@
 $1+\sum_{m=1}^{\infty}\frac{q^{m(m-1)}}{(q^{2};q^{2})_{m}}z^{m}=\prod_{m=0}^{\infty}(1+zq^{2m}).$
-Hence
+Hence it holds that
 $a_{n}$ $=$ $\frac{(-q)^{n}}{(q^{2};q^{2})_{\infty}}\sum_{m=0}^{\infty}\frac{q^{m(m-1)}}{(q^{2};q^{2})_{m}}(-q^{2(n+1)})^{m}$
@@
 which concludes the proof of Theorem 3. ∎
-It can be shown that $V_{\lambda}(\varphi_{\mathrm{ortho}})=V_{\lambda}(\varphi)$, where $V_{\lambda}(\varphi)$ is again the shift-invariant space (2) with Gaussian generating function $\varphi$ as given in (9).
-In Fig.
-2, the function $\varphi_{\mathrm{ortho}}$ together with some of its translates is depicted for bandwidth parameter $\beta=100$ and $\lambda=3/\beta$.
+In Fig. 2, the function $\varphi_{\mathrm{ortho}}$ along with some of its translates is depicted for bandwidth parameter $\beta=100$ and $\lambda=3/\beta$.
+Proposition 1
+The function $\phi$ in (21) has for $x\in\mathbb{R}$ the representation
+$\phi(x)=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{{(q^{2};q^{2})_{n}}}\,\varphi_{\mathrm{ortho}}(x-n\lambda),$ (24)
+where $q$ is as in Theorem 3
+Proof:
+The function $\varphi_{\mathrm{ortho}}$ as given in Eq. (20) has the Fourier transform
+$\hat{\varphi}_{\mathrm{ortho}}(\omega)$ $=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{-\mathrm{i}\omega x}\varphi_{\mathrm{ortho}}(x)\,\mathrm{d}x$
+$=\frac{1}{Q_{0}^{1/2}}\sum_{n=0}^{\infty}\frac{(-q)^{n}}{(q^{2};q^{2})_{n}}\,\mathrm{e}^{-\mathrm{i}n\lambda\omega}\hat{\phi}(\omega)$
+$=\frac{1}{Q_{0}^{1/2}}\sum_{n=0}^{\infty}\frac{(-q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}}{(q^{2};q^{2})_{n}}\,\hat{\phi}(\omega).$
+Because of Eq. (17) it holds that
+$\sum_{n=0}^{\infty}\frac{(-q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}}{(q^{2};q^{2})_{n}}=\prod_{n=0}^{\infty}(1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}\cdot q^{2n})^{-1}.$
+We thus obtain
+$\hat{\varphi}_{\mathrm{ortho}}=\frac{1}{Q_{0}^{1/2}}\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]^{-1}\hat{\phi}(\omega),$
+which is equivalent to
+$\hat{\phi}(\omega)=Q_{0}^{1/2}\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]\,\hat{\varphi}_{\mathrm{ortho}}(\omega).$ (25)
+Because of Eq. (16) it furthermore holds that
+$\prod_{n=0}^{\infty}[1+q\mathrm{e}^{-\mathrm{i}\lambda\omega}(q^{2})^{n}]$ $=\sum_{n=0}^{\infty}\frac{q^{n(n-1)}}{(q^{2};q^{2})_{n}}(q\mathrm{e}^{-\mathrm{i}\lambda\omega})^{n}$
+$=\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{(q^{2};q^{2})_{n}}\,\mathrm{e}^{-\mathrm{i}n\lambda\omega}.$
+Inserting the last expression into Eq. (25) and inverting the Fourier transform yields
+$\phi(x)$ $=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{(q^{2};q^{2})_{n}}\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{e}^{\mathrm{i}\omega(x-n\lambda)}\hat{\varphi}_{\mathrm{ortho}}(\omega)\,\mathrm{d}\omega$
+$=Q_{0}^{1/2}\sum_{n=0}^{\infty}\frac{q^{n^{2}}}{{(q^{2};q^{2})_{n}}}\,\varphi_{\mathrm{ortho}}(x-n\lambda),$
+which concludes the proof of Proposition 1. ∎
+Corollary 1
+Let the Gaussian generator $\varphi$ be as given in (9) and let the function $\varphi_{\mathrm{ortho}}$ be as given in Eq. (20). Then for any $\lambda>0$ it holds that
+$V_{\lambda}(\varphi_{\mathrm{ortho}})=V_{\lambda}(\varphi).$ (26)
+Proof:
+For any generating function $\varphi\in L^{2}(\mathbb{R})$ and any $\lambda>0$, the shift-invariant space $V_{\lambda}(\varphi)$ as given in (2) may also be defined as the closed linear span of the system of functions $\{\varphi(\cdot-n\lambda);\,n\in\mathbb{Z}\}$ in $L^{2}(\mathbb{R})$; in particular, $V_{\lambda}(\varphi)$ is a closed linear subspace of $L^{2}(\mathbb{R})$.
+In the case of the above Gaussian generator $\varphi$, due to Theorem 3, $\varphi_{\mathrm{ortho}}\in V_{\lambda}(\phi)=V_{\lambda}(\varphi)$ so that
+$V_{\lambda}(\varphi_{\mathrm{ortho}})\subseteq V_{\lambda}(\varphi).$ (27)
+On the other hand, by reason of Proposition 1, $\varphi\propto\phi\in V_{\lambda}(\varphi_{\mathrm{ortho}})$ so that
+$V_{\lambda}(\varphi)\subseteq V_{\lambda}(\varphi_{\mathrm{ortho}}).$ (28)
+The set inclusion (27) in combination with the set inclusion (28) results in the set equality (26), which concludes the proof of Corollary 1. ∎
 IV Applications
@@
 $\sum_{n\in\mathbb{Z}}\Phi_{\mathrm{int}}(n\lambda)\mathrm{e}^{-\mathrm{i}n\lambda\omega}=\frac{\Lambda}{\sqrt{2\pi}}\sum_{k\in\mathbb{Z}}\hat{\Phi}_{\mathrm{int}}(\omega+k\Lambda)=1$ (29)
-(the left-hand equation being Poisson’s summation formula; e.g., [1]), which is equivalent to $\Phi_{\mathrm{int}}(n\lambda)=\delta_{n},\,n\in\mathbb{Z}$.
+(the first equation being Poisson’s summation formula; e.g., [1]), which is equivalent to $\Phi_{\mathrm{int}}(n\lambda)=\delta_{n},\,n\in\mathbb{Z}$.
 Therefore, the set of shifted pulses $\{\Phi_{\mathrm{int}}(x-n\lambda);n\in\mathbb{Z}\}$ is ISI-free at points in time $x_{n}=n\lambda$.
-The right-hand equation in (23) is, of course, the well-known Nyquist criterion for ISI-free pulses [10].
-Concerning the use of the also ISI-free pulses generated by (15) (for arbitrary $\beta>0,\,\lambda>0$) refer to [11].
+The second equation in (29) is, of course, an instance of the well-known Nyquist criterion for ISI-free pulses [10].
+Concerning the use of the also ISI-free pulses generated by (15) (for arbitrary positive parameters $\beta$ and $\lambda$) refer to [11].
 IV-B Orthonormal Pulses with ISI-Free Matched Filter Output
 IV-B1 Orthonormal Pulses
-It is well-known [1] that for an arbitrary function $\varphi\in L^{2}(\mathbb{R})$ the system of functions $\{\varphi(x-n\lambda);n\in\mathbb{Z}\}$ forms an orthonomal system in $L^{2}(\mathbb{R})$ if and only if in (8) $A=B=1$ may be chosen.
-For the function $\varphi_{\mathrm{ortho}}$ of Theorem 3 this is true due to Eq.
-(21) and the right-hand equation in (23).
+For any function $\varphi\in L^{2}(\mathbb{R})$ the system of functions $\{\varphi(x-n\lambda);n\in\mathbb{Z}\}$ forms an orthonormal system in $L^{2}(\mathbb{R})$ if and only if in (8) $A=B=1$ may be chosen [1].
+For the function $\varphi_{\mathrm{ortho}}$ of Theorem 3 this is true because of Eq.
+(22) and the second equation in (29).
 Therefore, $\varphi_{\mathrm{ortho}}$ may be used as a pulse shape to generate a set $\{\varphi_{\mathrm{ortho}}(x-n\lambda);\,n\in\mathbb{Z}\}$ of orthonormal pulses.
@@
 At the receiver, the filter $\boldsymbol{P}_{\varphi_{\mathrm{ortho}}}$ obtained by replacing the prefilter function $\varphi$ in (5) with $\varphi_{\mathrm{ortho}}$ forms a matched filter allowing optimal detection of the pulses $\varphi_{\mathrm{ortho}}(x-n\lambda)$ at points in time $x_{n}=n\lambda,\;n\in\mathbb{Z},$ in the presence of noise.
-Thanks to orthogonality, overlap of consecutive pulses is irrelevant.
-Moreover, since [cf.
+Due to orthogonality, overlap of adjacent pulses is of no relevance.
+Moreover, since it holds that [cf.
 (6) and Eq.
@@
 In the light of the last application, the two proposed pulse shapes $\Phi_{\mathrm{int}}$ and $\varphi_{\mathrm{ortho}}$ are seen to correspond to the classical pulse shapes produced by a raised-cosine filter or a root raised-cosine filter respectively [10].
-Recall that the proposed pulse shapes decay (super-)exponentially as $x\rightarrow\pm\infty$ whereas the classical pulse shapes merely decay as $x^{-3}$.
+Recall that the proposed pulse shapes decay (super-)exponentially to zero as $x\rightarrow\pm\infty$ whereas the classical pulse shapes merely decay order of $x^{-3}$.
 acknowledgement
-The author wishes to thank Ayush Bhandari for interesting discussions and helpful suggestions concerning some topics of this paper.
+The author wishes to thank Ayush Bhandari for interesting discussions and helpful suggestions concerning an earlier draft of this paper.